OpenStudy (anonymous):

5x-4y+12=0 find slope, help plz

OpenStudy (agent0smith):

Can you get it to look like y = ... (ie y = mx + b)

OpenStudy (anonymous):

4=5x+12 ??

OpenStudy (agent0smith):

No, you can't just move everything over like that... and where did the y go? First, subtract 12 on both sides of 5x-4y+12=0

OpenStudy (anonymous):

5x-4y=12-12=0-12?

OpenStudy (agent0smith):

I can't tell exactly what you're doing because you have an extra equals sign in there \[\Large 5x - 4y + 12 -12 = 0 -12\] which simplifies to what?

OpenStudy (anonymous):

5x-4y= -12 then divide by 5 to get x alone?

OpenStudy (anonymous):

or divide by 4 to get rid of it, so we have 5x= -3 ???

OpenStudy (agent0smith):

Hold on, you can't just magically make variables disappear like that just by dividing by 4. Get y alone BEFORE doing any dividing. First you have to subtract 5x on both sides of \[\Large 5x-4y= -12 \]

OpenStudy (anonymous):

with all do respect, If I knew how to do it. I would not be on this site asking for someone to show me. thanks for trying to help me but I'll just wait for someone else. have a good day

OpenStudy (agent0smith):

I am helping you, i'm correcting the errors you're making and teaching you the correct steps.

OpenStudy (anonymous):

I understand, I just don't have the time to sit here and constantly be corrected. I would just go to my professors office hours if I wanted that. Thanks anyways. have a good day

OpenStudy (agent0smith):

So you'd rather keep making mistakes then, and never be corrected... why did you come here then? Just for an answer? What you said doesn't make sense. Noticing and correcting your errors is part of how you learn.

OpenStudy (agent0smith):

What exactly were you hoping for? Someone to just show the steps and that's it?

OpenStudy (anonymous):

Yes, she was asking for the steps on how to do it. Then she could of done it and solved it! Girls hate to be corrected. trust me

OpenStudy (agent0smith):

Yeah... but there's no point knowing the right steps to something if you also are unaware what steps are wrong, since knowing the right steps won't prevent you taking a wrong step occasionally. I see smart math kids, who know the right steps, occasionally make a mistake just because they didn't know it was incorrect. And while i'm here... "with all do respect"... it's "with all DUE respect".

OpenStudy (anonymous):

Yeah, you are a jerk!